OpenStudy (anonymous):

Find the exact circumference, in terms of , of a circle with the given radius. miles

OpenStudy (anonymous):

x/2 miles

OpenStudy (cwrw238):

use the formula C = 2 * pi * r where r = x/2

OpenStudy (anonymous):

thank you! thank you! thank you!!

OpenStudy (cwrw238):

no probs